Write the letter corresponding to the slope of the line passing through the points (–8, 8) and (7, 8).
A) 15

B) 0

C) Undefined

D) 1


Answer by Nate(3500)

Write the letter corresponding to the slope of the line passing through the points (–8, 8) and (7, 8).
slope = (y-y)/(x-x)
slope = (8-8)/(-8-7) = 0/(-15) = 0
Your solution should be (B).